what volume of a 0.50 M KBr solution is needed to provide 3.0 moles of KBr
Answer #1

Ans. Note that molarity = moles of solute per liter of solution.

Let the required volume of solution is X liters.

Now,

            Number of moles = Molarity x Volume in liters

            Or, 3.0 moles = 0.50 M x X L

            Or, 3.0 moles = 0.50 mol L-1 x X L = 0.50X mol

            Hence, X = 3.0 / 0.50 = 6.00

Therefore, required volume of solution = 6.00 liters